OpenStudy (anonymous):

Find the prime factorization of 840. Please show all of your work

OpenStudy (anonymous):

840/2=420 420/2=210 210/2=105 105/3=35 35/5=7 So 2^3*3*5*7

OpenStudy (anonymous):

GR8 THANX BUD